# Ergodic Theorems and Equivalence of Green’s Kernel for Random Walks in Random Environments

### v3

@@ 8.1. Action of Linear Transformations:
Theorem 3.7 holds under its assumptions if the local process is stationary under some measure $\mathbb{Q}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ which is mutually absolutely continuous to $\mu$.
We shall develop in the next section the proof of Lemma 8.5.
@@
The first part of the proof is to construct $\mathbb{Q}$.
Let $\lambda^{\prime}$ be an invariant measure of $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 0}$. First, we make an observation on this measure.
Let $\alpha:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ given by $\alpha(e_{i})=-e_{i}\quad i=1,2,\dots,d$.
Lemma 9.1.
$\lambda^{\prime}$ is an $\alpha$-invariant measure. More precisely,
$\lambda^{\prime}\circ\alpha(B)=\lambda^{\prime}(B)\quad\forall B\in\mathcal{B}(d)^{\prime}.$
Proof.
Start by noting that for every $B\in\mathcal{B}(d)^{\prime}$,
$\int_{\mathcal{S}(d)}K_{d}^{\Lambda}(\omega,B)d\lambda^{\prime}\circ\alpha=\int_{\mathcal{S}(d)}K_{d}^{\Lambda}(\omega,\alpha(B))d\lambda^{\prime}=\lambda^{\prime}\circ\alpha(B).$
This shows $\lambda^{\prime}\circ\alpha$ is also an invariant stationary measure for the environment process. Furthermore $\lambda^{\prime}\circ\alpha$ is mutually absolutely continuous to $\mu\circ\alpha$. One can also show that $\mu\circ\alpha$ is stationary and ergodic with respect to the group-of-shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$.
Let $g:\mathcal{S}(d)\to\mathbb{R}$ be a measurable function. By Birkhoff’s Ergodic Theorem,
$\frac{1}{n}\sum_{i=0}^{n-1}g(\tau_{X_{\Lambda(\omega)}(n)}\omega)$ $\to\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\quad\omega\penalty 10000\ \mu\textit{-a.e.}$ (7)
$\frac{1}{n}\sum_{i=0}^{n-1}g(\tau_{X_{\Lambda(\omega)}(n)}\omega)$ $\to\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\circ\alpha\quad\omega\penalty 10000\ \mu\circ\alpha\textit{-a.e.}$ (8)
Therefore, we conclude that,
$\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}=\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\circ\alpha,$
which completes the proof. ∎
Fix $\omega\in\mathcal{S}(d)$.
@@
$\mathbb{P}[X_{\omega}(1)=x+v\Big|X_{\omega}(0)=x]=\omega(x,v)\quad v\in V_{d}.$ (9)
@@
$\mathbb{P}[X_{\alpha(\omega)}(1)=x+v\Big|X_{\alpha(\omega)}(0)=x]=\omega(x,-v).$ (10)
@@
where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ kernels of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk. Further $\{\gamma_{n}\}_{n\geq 1}$ is a sequence of i.i.d. Bernoulli random variables.
We therefore define the conditional environment process
@@
This is the control argument using Bernoulli variables as we had mentioned, and now we shall use it explicitly to demonstrate the stationarity of the local process.
We claim that $\mathbb{Q}=\lambda^{\prime}$.
@@
Let $n\geq 1$. Let $B_{0},B_{1},\dots,B_{n}\in\mathcal{B}(d)^{\prime}$. Let the probability measure associated with $\omega$ be $\lambda^{\prime}$. Let $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$. Then for every $1\leq k\leq n-1$, we have
@@
$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\ \bar{\omega}(k+1)\in\alpha(B_{k+1}),\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,1-\gamma_{k+1},\ldots,\gamma_{n}\Bigr].$ (11)
For $k=0$,
$\lambda^{\prime}\Big[\omega\in\alpha(B_{0}),\,\bar{\omega}(1)\in B_{1},\,\ldots,\,\bar{\omega}(k)\in B_{k},\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,\gamma_{n}\Bigr]$
$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in\alpha(B_{1})\ldots,\,\bar{\omega}(k)\in B_{k}\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,1-\gamma_{1},\ldots,\gamma_{k},\ldots,\gamma_{n}\Bigr].$ (12)
@@
The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{1}),\dots,\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$. Now, conclude stationarity using the stationarity of $\bar{\omega}(n)$ and taking $\gamma_{i}=1$. It is clear that
@@
By Lemma 9.1 and Lemma 9.2, the integral reduces to
$\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\lambda^{\prime}(\alpha(\omega)\in B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma).$
Suppose the result is true for $n=m$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity, as others follow from the technique of proof. For $k=m+1$, we write
@@
This reduces by definition to the desired form, which completes the proof.

### v4

@@ 8.1. Action of Linear Transformations:
Let $\mathbb{Q}$ be a measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic. Then $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}=\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ in distribution (and therefore both are stationary sequences) under $\mathbb{Q}$.
@@
Let $\lambda^{\prime}$ be the stationary ergodic measure of the chain $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 0}$. We set $\mathbb{Q}=\lambda^{\prime}$. Fix $\omega\in\mathscr{S}(d)$.
@@
(6) $\mathbb{P}[X_{\omega}(1)=x+v\Big|X_{\omega}(0)=x]=\omega(x,v)\quad v\in V_{d}.$
@@
(7) $\mathbb{P}[X_{\alpha(\omega)}(1)=x+v\Big|X_{\alpha(\omega)}(0)=x]=\omega(x,-v).$
@@
where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ kernels of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk. Further $\{\gamma_{n}\}_{n\geq 1}$ is a sequence of i.i.d. Bernoulli random variables. We therefore define the conditional environment process
@@
This is the control argument using Bernoulli variables as we had mentioned, and now we shall use it explicitly to demonstrate the stationarity of the local process or the equivalent statement of Lemma 8.5. Let $\bar{\omega}(k)$ denote the environment process $\tau_{X_{\Lambda(\omega)}(k)}\omega$ for $k\geq 0$.
@@
Let $n\geq 1$. Let $B_{0},B_{1},\dots,B_{n}\in\mathcal{B}(d)^{\prime}$. Let the probability measure associated with $\omega$ be $\lambda^{\prime}$. Then for every $0\leq k\leq n-1$, we have
@@
$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\ \bar{\omega}(k+1)\in\alpha(B_{k+1}),\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,1-\gamma_{k+1},\ldots,\gamma_{n}\Bigr].$
@@
The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),\dots,\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$. Since $\{\bar{\omega}(k)\}_{k\geq 0}$ is stationary under $\lambda^{\prime}$, this establishes the assertion of Lemma 8.5 and therefore stationarity of $\{\pi_{0}(\tau_{X_{\omega}(k)}\omega)\}_{k\geq 0}$. Note that taking $\gamma_{i}=1$, it is clear that
@@
This shows that choosing $\gamma_{i}$’s controls the path suitably.
@@
By Lemma 9.1, the integral reduces to
$\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\mathbb{E}[\boldsymbol{1}_{\omega}(\alpha(B_{0}))\mathbb{E}[\boldsymbol{1}_{\bar{\omega}(1)}(\alpha.B_{1})\Big|\omega,1-\gamma]]$
This reduces to $\lambda^{\prime}[\omega\in\alpha.B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma]$.
Suppose the result is true for $n=m\geq 1$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity, as others follow from the technique of proof. For $k=m+1$, we write
@@
This reduces by definition to the desired form, which completes the proof of Lemma 9.2.